One kind of e-book display consists of millions of very small spheres that float in a thin fluid layer between two conducting, transparent plates. Each sphere is black on one side and white on the other, and possesses an electric dipole moment directed from the white side to the black. When an electric field is applied between the plates, the spheres rotate so that their dipole moment lines up with the field. Depending on the field's direction, either the black or the white sides of the spheres can be made visible.
The dipole moment of the spheres can be modeled as two opposite charges of magnitude 3.8×10−15 C separated by a distance of 100 μm.

You may want to review (Pages 654 - 655) .

Part A

What is the maximum possible torque on a sphere if the electric field between the transparent plates is 4.9×105 N/C ?

Express your answer to two significant figures and include the appropriate units.

τ

τ

=

nothingnothing

SubmitRequest Answer
Answer #1

Torque on dipole in electric field is gvien by:

Torque = pxE = p*E*sin theta

p = dipole moment = q*d

E = electric field = 4.9*10^5 n?C

For max torque, sin theta = 1

So,

Max torque = q*d*E

Using given values:

Max torque = 3.8*10^-15*100*10^-6*4.9*10^5

Ma torque = 1.9*10^-13 N-m

Please Upvote.